In the following table is a partial computer output based on a sample of 21 observations, relating an independent variable (X) and a dependent variable:

Predictor    Coefficient Standard Error
Constant 30.139 1.181
X ‐0.2520 0.022
SOURCE    SS
Regression 1,759.481
Error    259.186

a. Develop the estimated regression line.

b. At α = 0.05, test for the significance of the slope.

c. At α = 0.05, perform an F‐test.  

d. Determine the coefficient of determination.

e. Determine the coefficient of correlation.
Answer #1

a) Estimated regression line is given by

hat{Y} = b_0 + b_1hat{X}

where bo,bi are the estimated coefficients of original coefficients eta_0,eta_1 and hat{Y} is the fitted value value.

Thus, y 30.139-0.2520x

b)

  • Hypothesis: Ho: The slope of the regression line is equal to zero.

Ha: The slope of the regression line is not equal to zero.

If the relationship between Y and X is significant, the slope will not equal zero.

  • Given the significance level is 0.05, we will conduct a t-test using sample data.
  • To apply the t-test to sample data, we require the standard error of the slope, the slope of the regression line, the degrees of freedom, the t statistic test statistic, and the P-value of the test statistic.

    From the table, we can see that-

    b1 = -0.2520 SE = 0.022

    We compute the degrees of freedom and the t statistic test statistic, using the following equations.

    DF = n - 2 = 21 - 2 = 19

    t = b1/SE = -0.2520/0.022 = -11.45

  • Based on the t statistic test statistic and the degrees of freedom, we determine the P-value. The P-value is the probability that a t statistic having 19 degrees of freedom is more extreme than -11.45. Since this is a two-tailed test, using the table, we find that at 5% significance level, P-value < 0.0001

  • Since the P-value is less than the significance level (0.05), we cannot accept the null hypothesis and the result is significant.

c)

To test the hypothesis Ho 61 0, the statistic used is based on the b13be40341fd44a1f4b363313147aeb7.png distribution. It can be shown that if the null hypothesis ece6f3083be382dd642864fca5b1696f.png is true, then the statistic:

MSR SSR/1 MS SSE/ (n 2) Fo


follows the b13be40341fd44a1f4b363313147aeb7.png distribution with c50e41b438d145094410c714ab03f00c.png degree of freedom in the numerator and (n-2) degrees of freedom in the denominator. ece6f3083be382dd642864fca5b1696f.png is rejected if the calculated statistic, Fo, is such that:

c1b43a2d045812dad1c733232dedb6d6.png


where Ja,1,n-2 is the percentile of the b13be40341fd44a1f4b363313147aeb7.png distribution corresponding to a cumulative probability of (22b1b50a81911a480a51fa965d89c423.png) and 4bc6c42bbabe567d1f2516326e52b775.png is the significance level.

From the table, we can see that regression sum of squares

SS_R = sum_{i=1}^{n} (hat{y_i} - ar{y})^2 = 1759.481

The error sum of squares can be calculated as:

SSE = Σ(Vi-Vi) (Viーの一259.186 i= 1 259.186

Knowing the sum of squares, the statistic to test Ho 61 0 can be calculated as follows:

fo= MSR = MSE SSR/1 SSE/(n-2)

1759.481/1759.481_128.985 259.186/(19) 13.641

The critical value at a significance level of 0.05 is J0.05,1,19 = 4.381. Since 4fa4e9e767b6970b3a40d5e6e0df41fe.png, Ho 61 0 is rejected and it is concluded that 3f912a9f94bb8705de28c11fb1dcb786.png is not zero, implying that a relation does exist between Y and X for the data in the preceding table.

d)

The coefficient of determination is a measure of the amount of variability in the data accounted for by the regression model.

The coefficient of determination is the ratio of the regression sum of squares to the total sum of squares.

5 SSR SST
90bd02aebd582db34293c450956157b9.png can take on values between 0 and 1 .

SSr-〉·(Vi _ у)2ーSSR + SSE 1759.481 + 259.186 = 2018.667

  5 SSR SST= rac{1759.481}{2018.667} = 0.872